# A note on generation and descent for derived categories of noncommutative schemes

### v3

@@ 3.2. Results
If $E\in\operatorname{coh}\mathcal{A}$, then $\pi_{\ast}E\in\operatorname{coh}X$ by Remark 3.1.1. This establishes a well-defined functor $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$. To verify exactness, consider a short exact sequence in $\operatorname{coh}\mathcal{A}$:
@@
Lemma 3.2.2.
Let $X$ be a Noetherian scheme. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$, then there exists an exact functor $\pi_{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(X)$ of triangulated categories.
@@
Lemma 3.2.1 tells us $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$ is exact on abelian categories, so there exists an induced functor on derived categories $D_{\operatorname{coh}}(\mathcal{A})\to D_{\operatorname{coh}}(X)$. To see $\pi_{\ast}$ preserves bounded cohomology, this follows from the fact that $\pi$ is exact. ∎
Remark 3.2.3.
Let $R$ be a commutative Noetherian ring. Suppose $P,Q\in\operatorname{perf}R$. The following are equivalent: $P$ finitely builds $Q$ if, and only if, $\operatorname{Supp}(Q)\subseteq\operatorname{Supp}(P)$. This is [Nee92, Lemma 1.2] and [Hop87, Theorem 11].
Remark 3.2.4.
Let $X$ be a Noetherian scheme. Suppose that $j\colon U\to X$ is an open subscheme. Choose $E,P\in D^{b}_{\operatorname{coh}}(X)$. If $\mathbb{L}j^{\ast}E\in\langle\mathbb{L}j^{\ast}P\rangle$ in $D^{b}_{\operatorname{coh}}(U)$, then there exists $C\in D^{b}_{\operatorname{coh},Z}(X)$ such that $E\in\langle P\oplus C\rangle$ in $D^{b}_{\operatorname{coh}}(X)$ where $Z=X\setminus U$. This is [Lan23, Lemma 4.1.2].
Remark 3.2.5.
Let $X$ be a Noetherian scheme. If $X=\bigcup^{n}_{i=1}Z_{i}$ denotes the maximal irreducible components and $G_{i}\in D^{b}_{\operatorname{coh}}(Z_{i})$ is a classical generator for each $1\leq i\leq n$, then $\oplus_{i=1}^{n}\mathbb{R}\pi_{i,\ast}G_{i}$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$ where $\pi_{i}\colon Z_{i}\to X$ is the closed immersion. This can be shown from [DL24, Proposition 3.3].
Proof of Theorem 3.1.
If $X=\emptyset$, there is nothing to check, so without loss of generality, $X$ is non-empty. Note any closed subscheme $Z$ of $X$ is Noetherian $J\textrm{-}2$. By Noetherian induction, we assume the claim holds for any properly contained closed subscheme $Z$ of $X$.
First, consider the case where $X$ is an integral scheme. Then $\operatorname{reg}(X)$ is a nonempty open subscheme of $X$. Let $i\colon U\to X$ be an open immersion, with $U$ an open affine subscheme contained in $\operatorname{reg}(X)$. Fix a classical generator $E$ for $D^{b}_{\operatorname{coh}}(X)$. Then $i^{\ast}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(U)$ by Remark 3.2.3. Thus, $i^{\ast}\pi_{\ast}G$ finitely builds $i^{\ast}E$ in $D^{b}_{\operatorname{coh}}(U)$. Consequently, there exists $A\in D^{b}_{\operatorname{coh}}(Z)$, where $Z$ is a closed subscheme contained in $X\setminus U$, with a closed immersion $j\colon Z\to X$, such that $E$ is finitely built by $\pi_{\ast}G\oplus j_{\ast}A$; cf. Remark 3.2.4. We then demonstrate that $\pi_{\ast}G$ finitely builds $j_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$. There exists a commutative diagram of $\mathcal{O}_{X}$-algebras:
${{\mathcal{A}}}$${{j^{\ast}\mathcal{A}}}$${{\mathcal{O}_{X}}}$${{j_{\ast}\mathcal{O}_{Z}}.}$$\scriptstyle{\phi_{1}}$$\scriptstyle{\phi_{2}}$$\scriptstyle{\theta}$$\scriptstyle{\pi}$
Since $Z$ is a properly contained subscheme of $X$, the induction hypothesis states that if $G^{\prime}$ is a classical generator for $D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})$, then $\theta_{\ast}G^{\prime}$ is a classical generator for $D^{b}_{\operatorname{coh}}(Z)$. Consequently, $(\phi_{2}\circ\theta)_{\ast}G^{\prime}$ finitely builds $j_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$. Moreover, since $G$ finitely builds $\phi_{1,\ast}G^{\prime}$ in $D^{b}_{\operatorname{coh}}(\mathcal{A})$, we have that $\pi_{\ast}G$ finitely builds $(\pi\circ\phi_{1})_{\ast}G^{\prime}$ in $D^{b}_{\operatorname{coh}}(X)$. Hence, $\pi_{\ast}G$ finitely builds $j_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$ and, consequently, finitely builds $E$.
Lastly, in the general case where $X$ is not an integral domain, a similar argument applies to show that $\pi_{\ast}G$ finitely builds any classical generator for each maximal irreducible component of $X$, cf. Remark 3.2.5. ∎
Remark 3.2.6.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$. This is [DL24, Corollary 3.6].
Corollary 3.2.7.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian $J\textrm{-}2$ schemes of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{Y}$-algebra with structure map $\eta\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}(\pi\circ\eta)_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
This is Remark 3.2.6 and Theorem 3.1. ∎
Lemma 3.2.8.
Let $R,S$ be Noetherian rings and $\pi\colon R\to S$ a ring homomorphism. If $E\in D{\operatorname{Qcoh}}(R)$ and $A\in D{\operatorname{Qcoh}}(S)$ is a complex of $(S,S)$-bimodules, then there exists a isomorphism in $D{\operatorname{Qcoh}}(R)$:
$E\overset{\mathbb{L}}{\otimes}_{R}\pi_{\ast}A\cong\pi_{\ast}(\mathbb{L}\pi^{\ast}E\overset{\mathbb{L}}{\otimes}_{S}A).$
@@
We can realize $\phi_{\ast}$ and $\mathbb{L}\phi^{\ast}$ as the derived tensor product functors:
@@
where $S_{R}$ denotes $S$ as a right $R$-module via $\phi$. There exists an isomorphism in $D{\operatorname{Qcoh}}(R)$:
@@
Before obtaining our desired results, we recall a useful statement relating generation time to the number of cones needed to finitely build perfect complexes. The following result is [LO24, Theorem 1.2] which significantly sharpens Remark 2.10.
Theorem 3.2.9 (Lank-Olander).
If $X$ is a Noetherian scheme, $G$ is an object of $D^{b}_{\operatorname{coh}}(X)$, $d$ a non-negative integer, then the following are equivalent:
  1. (1)
    $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{d}$
  2. (2)
    $\operatorname{perf}X\subseteq\langle G\rangle_{d}$.
From this result, we now are able to prove Corollary 3.2.
Proof of Corollary 3.2.
If $D^{b}_{\operatorname{coh}}(S)$ has infinite Rouquier dimension, then there is nothing to check, so without loss of generality we may assume this value is finite. Suppose $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(S)$ with generation time is $g$. By Theorem 3.1, $\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(R)$. Choose $n\geq 0$ such that $R\in\langle\pi_{\ast}G\rangle_{n}$. According to Lemma 3.2.8, if $P\in\operatorname{perf}R$, the projection formula ensures $P\in\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)\rangle_{n}$. In particular, this tells us that $\operatorname{perf}R\subseteq\langle\pi_{\ast}G\rangle_{n(g+1)}$. Now, we will utilize this fact in the following. However, the remainder of follows by Theorem 3.2.9. ∎
Remark 3.2.10.
There exists a similar result if, a priori, there exists $E\in D^{b}_{\operatorname{coh}}(S)$ which is a complex of $(S,S)$-bimodules such that $\pi_{\ast}E$ finitely builds $R$ in $D^{b}_{\operatorname{coh}}(R)$. It is worthwhile to note that a similar strategy for showing Corollary 3.2 can be obtained via the proof of [OŠ12b, Theorem 7].
Remark 3.2.11.
Let $X$ be an embeddable variety over a field $k$. Suppose $\mathcal{A}$ is a sheaf of Azumaya algebras on $X$ with structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $E\in D^{-}(\mathcal{A})$ and $G\in D^{-}(X)$, then there exists an isomorphism in $D_{\operatorname{Qcoh}}(X)$:
$\pi_{\ast}(E\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}G)\cong\pi_{\ast}E\overset{\mathbb{L}}{\otimes}G.$
This is [Kuz06, Lemma 2.12].
Proof of Corollary 3.3.
If $D^{b}_{\operatorname{coh}}(\mathcal{A})$ has infinite Rouquier dimension, then there is nothing to show; assume this value is finite. Theorem 3.1 ensures that $\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. Let $P\in\operatorname{perf}X$. There exists $n\geq 0$ such that $\mathcal{O}_{X}\in\langle\pi_{\ast}G\rangle_{n}$. By Remark 3.2.11, we observe that
$P$ $\in\langle\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P\rangle_{n}$
$=\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)\rangle_{n}.$
This implies $\operatorname{perf}X\subseteq\langle\pi_{\ast}G\rangle_{(g+1)n}$, where $g$ is the generation time. The desired claim follows by Theorem 3.2.9. ∎
Remark 3.2.12.
Let us keep the notation of Corollary 3.3. The following observation provides a weaker upper bound on Rouquier dimesion, but its techniques are worthwhile to write out. If $D^{b}_{\operatorname{coh}}(\mathcal{A})$ has infinite Rouquier dimension, then there is nothing to show; assume this value is finite. Theorem 3.1 ensures that $\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. Let $P\in\operatorname{perf}X$. There exists $n\geq 0$ such that $\mathcal{O}_{X}\in\langle\pi_{\ast}G\rangle_{n}$. By Remark 3.2.11, we observe that
$P$ $\in\langle\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P\rangle_{n}$
$=\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)\rangle_{n}.$
This implies $\operatorname{perf}X\subseteq\langle\pi_{\ast}G\rangle_{(g+1)n}$, where $g$ is the generation time. Hence, Remark 2.9 ensures $E\in\langle\pi_{\ast}G\rangle_{2(g+1)n}$, and the claim follows.
Proposition 3.2.13.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme, and $\mathcal{A}$ be a coherent $\mathcal{O}_{X}$-algebra. If there exists an open affine immersion $i\colon U\to X$ such that $\mathcal{O}_{U}\in\langle i^{\ast}\mathcal{A}\rangle_{1}$ and $U$ is nonempty, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least that of $D^{b}_{\operatorname{coh}}(U)$.
@@
Let $i\colon U\to X$ be an open immersion where $U=\operatorname{Spec}(R)$ is a nonempty affine scheme. Consider the Verdier localization $i^{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$. This tells us that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least that of $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$. Since $U$ is affine and $\mathcal{O}_{U}\in\langle i^{\ast}\mathcal{A}\rangle_{1}$, Corollary 3.2 tells us that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$ is at least that of $D^{b}_{\operatorname{coh}}(U)$, and this completes the proof. ∎
Proposition 3.2.14.
Let $X$ be an irreducible, Jacobson, catenary, $J\textrm{-}2$, and Noetherian scheme of finite Krull dimension. If $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is at least the Krull dimension of $X$.

### v4

@@ 4. Descent along structure morphism
If $E\in\operatorname{coh}\mathcal{A}$, then $\pi_{\ast}E\in\operatorname{coh}X$ by Remark 3.1. This gives us a functor $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$. To verify exactness, consider a short exact sequence in $\operatorname{coh}\mathcal{A}$:
@@
Lemma 4.2.
Let $X$ be a Noetherian scheme and $\mathcal{A}$ be a coherent $\mathcal{O}_{X}$-algebra. Then the canonical map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$ induces a exact functor $\pi_{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(X)$.
@@
[…]
Remark 4.3.
Let $X$ be a Noetherian scheme. If $X=\bigcup^{n}_{i=1}Z_{i}$ denotes the maximal irreducible components and $G_{i}\in D^{b}_{\operatorname{coh}}(Z_{i})$ is a classical generator for each $1\leq i\leq n$, then $\oplus_{i=1}^{n}\mathbb{R}\pi_{i,\ast}G_{i}$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$ where $\pi_{i}\colon Z_{i}\to X$ is the closed immersion. This can be shown from [DL24b, Example 3.4].
Theorem 4.4.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support and canonical map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
Any closed subscheme $Z$ of $X$ is Noetherian $J\textrm{-}2$. We will prove the claim by Notherian induction on $X$. If $X=\emptyset$, there is nothing to check, so we can impose that $X$ is non-empty. Without loss of generality, we can assume the claim holds for any properly contained closed subscheme $Z$ of $X$.
Consider the case where $X$ is an integral scheme. Then $\operatorname{reg}(X)$ is a nonempty open subscheme of $X$ as it contains the generic point. Let $j\colon U\to X$ be an open immersion, with $U$ an affine open subscheme contained in $\operatorname{reg}(X)$. There is a a classical generator $E$ for $D^{b}_{\operatorname{coh}}(X)$ because $X$ is $J\textrm{-}2$, see [ELS20, Theorem 4.15]. Note that $j^{\ast}\pi_{\ast}G$ is a perfect complex on $U$ with full support because $U$ is regular and $\pi_{\ast}G$ has full support on $X$. This tells us that $j^{\ast}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(U)$, see [Nee92, Lemma 1.2]. Thus, $j^{\ast}\pi_{\ast}G$ finitely builds $j^{\ast}E$ in $D^{b}_{\operatorname{coh}}(U)$. Consequently, there exists $A\in D^{b}_{\operatorname{coh}}(Z)$, where $Z$ is a closed subscheme contained in $X\setminus U$, with a closed immersion $i\colon Z\to X$, such that $E$ is finitely built by $\pi_{\ast}G\oplus i_{\ast}A$, see [Lan24, Lemma 4.1.2].
Next, we demonstrate that $\pi_{\ast}G$ finitely builds $i_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$. Let $\theta\colon\mathcal{O}_{Z}\to i^{\ast}\mathcal{A}$ be the structure morphism. There exists a commutative diagram of exact functors [^6] This follows from Remark 3.3, [BDG17b, Proposition 3.12], and Lemma 4.2.:
${{D^{b}_{\operatorname{coh}}(Z)}}$${{D^{b}_{\operatorname{coh}}(X)}}$${{D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})}}$${{D^{b}_{\operatorname{coh}}(\mathcal{A})}.}$$\scriptstyle{i_{\ast}}$$\scriptstyle{i_{\ast}}$$\scriptstyle{\pi_{\ast}}$$\scriptstyle{\theta_{\ast}}$
Note that $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$ admits a classical because $Z$ is $J\textrm{-}2$ and $i^{\ast}\mathcal{A}$ is a coherent $\mathcal{O}_{Z}$-algebra, see [ELS20, Theorem 4.15]. Since $Z$ is a properly contained subscheme of $X$, the induction hypothesis ensures that if $G^{\prime}$ is a classical generator for $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$, then $\theta_{\ast}G^{\prime}$ is a classical generator for $D^{b}_{\operatorname{coh}}(Z)$. Consequently, $(i\circ\theta)_{\ast}G^{\prime}$ finitely builds $i_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$. Moreover, since $G$ finitely builds $i_{\ast}G^{\prime}$ in $D^{b}_{\operatorname{coh}}(\mathcal{A})$, we have that $\pi_{\ast}G$ finitely builds $(\pi\circ i)_{\ast}G^{\prime}$ in $D^{b}_{\operatorname{coh}}(X)$. Hence, $\pi_{\ast}G$ finitely builds $i_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$ and, consequently, finitely builds $E$.
Now we work in the general case where $X$ is not integral. It suffices to show that $i_{\ast}D^{b}_{\operatorname{coh}}(Z)$ is contained in $\langle\pi_{\ast}D^{b}_{\operatorname{coh}}(\mathcal{A})\rangle$ for each closed immersion $i\colon Z\to X$ from an irreducible component, see Remark 4.3. Let $\theta\colon\mathcal{O}_{Z}\to i^{\ast}\mathcal{A}$ be the structure morphism. Our work above tells us that $\langle\theta_{\ast}D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})\rangle=D^{b}_{\operatorname{coh}}(Z)$ because $Z$ is integral. However, we know that $i_{\ast}D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$ is contained in $D^{b}_{\operatorname{coh}}(\mathcal{A})$ (see Remark 3.7), and so, $i_{\ast}D^{b}_{\operatorname{coh}}(Z)$ belongs to $\langle\pi_{\ast}D^{b}_{\operatorname{coh}}(\mathcal{A})\rangle$ because $i_{\ast}\circ\theta_{\ast}=\pi_{\ast}\circ i_{\ast}$. This completes the proof. ∎
Corollary 4.5.
Let $f\colon Y\to X$ be a proper surjective morphism of Noetherian $J\textrm{-}2$ schemes of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{Y}$-algebra with canonical map $\pi\colon\mathcal{O}_{Y}\to\mathcal{A}$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}(f\circ\pi)_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
This is Theorem 4.4 coupled with [DL24b, Corollary 3.12]. ∎
Lemma 4.6.
Let $R,S$ be Noetherian rings where $R$ is commutative and $\phi\colon R\to S$ be a ring homomorphism. If $E\in D_{\operatorname{Qcoh}}(R)$ and $A\in D_{\operatorname{Qcoh}}(S)$ is a complex of $(S,S)$-bimodules, then there exists a isomorphism in $D_{\operatorname{Qcoh}}(R)$:
$E\overset{\mathbb{L}}{\otimes}_{R}\phi_{\ast}A\cong\phi_{\ast}(\mathbb{L}\phi^{\ast}E\overset{\mathbb{L}}{\otimes}_{S}A).$
Proof.
We can realize $\phi_{\ast}\colon D_{\operatorname{Qcoh}}(S)\to D_{\operatorname{Qcoh}}(R)$ and $\mathbb{L}\phi^{\ast}\colon D_{\operatorname{Qcoh}}(R)\to D_{\operatorname{Qcoh}}(S)$ as the derived tensor product functors [^7] These are respectively the derived functors for restriction and extension of scalars.:
@@
where $S_{R}$ denotes $S$ as a right $R$-module via $\phi$. There exists an isomorphism in $D_{\operatorname{Qcoh}}(R)$:
@@
Corollary 4.7.
@@
If $D^{b}_{\operatorname{coh}}(S)$ has infinite Rouquier dimension, then there is nothing to check, so without loss of generality we may assume this value is finite. Suppose $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(S)$ with generation time is $g$. By Theorem 4.4, $\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(R)$. Choose $n\geq 0$ such that $R\in\langle\pi_{\ast}G\rangle_{n}$. Let $P$ be in $\operatorname{perf}(R)$. Our hypothesis that $G$ is a complex of $(S,S)$-bimodules ensures that $\pi_{\ast}(\mathbb{L}\pi^{\ast}P\overset{\mathbb{L}}{\otimes}_{S}G)$ is isomorphic to $\pi_{\ast}G\overset{\mathbb{L}}{\otimes}_{R}P=P\overset{\mathbb{L}}{\otimes}_{R}S$, see Lemma 4.6. If we tensor with $P$ and use this isomorphism, then we see that $P\in\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}_{S}\mathbb{L}\pi^{\ast}P)\rangle_{n}$. In particular, this tells us that $\operatorname{perf}R\subseteq\langle\pi_{\ast}G\rangle_{n(g+1)}$. The desired claim follows from [LO24, Theorem 1.1] (as well as [OŠ12, Theorem 7]). ∎
Theorem 4.8.
Let $X$ be a scheme of finite Krull dimension which is integral, Jacobson [^8] See [Sta23, Tag 01P1]., catenary [^9] See [Sta23, Tag 02IV]., $J\textrm{-}2$, and Noetherian. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least the Krull dimension of $X$.